OpenStudy (anonymous):

sin(3A)/(sin(2A)-sin(A) = cos (??)

OpenStudy (anonymous):

express the first formula with only "cos"

OpenStudy (anonymous):

sin(3A) = sin(2A+A) (this is just the top) = sin(2A)cos(A)+sin(A)cos(2A) (trig identity of sin(a+b)) = sin(A+A)cos(A)+sin(A)cos(A+A) (sin(2A) = sin (A+A)) = 2sin(A)cos(A)cos(A)+sin(A)[cos(A)^2-sin(A)^2] (trig identity of cos(a+b) = 2sin(A)cos(A)^2 + sin(A)[(cos(A)^2 - (1- cos(A)^2)] (pythagorean identity sin^2+cos^2=1) = sin(A) [2cos(A)^2 + [cos(A)^2 - (1-cos(A)^2)] (factoring out a sin(A)) = sin(A) [4cos(A)^2 - 1] (adding like terms) = sin(A) [2cos(A) + 1][2cos(A) -1] (Factoring apart the 4cos(A)^2 - 1) So the top ends up being sin(A)[2cos(A) + 1][2cos(A) - 1] Now the bottom: Sin(2A) - Sin(A) = 2sin(A)cos(A) - sin(A) = sin(A)[2cos(A) - 1] notice we have: sin(A) and [2cos(A) - 1] on the top and bottom: sin(A)[2cos(A) + 1][2cos(A) - 1] / sin(A)[2cos(A) - 1] = 2cos(A) + 1

OpenStudy (anonymous):

wow thanks a lot !!